Please solve {{4/5x = -14}} showing all work and steps
Thank You

Answer by stanbon(75887)

4/5x = -14
---
Multiply both sides by (5/4) to get:
(5/4)(4/5)x = (5/4)(-14)
x = -70/4
x = -35/2
x = -17.5
=================
Cheers,
Stan H.